OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x))= x. f(x)= 4/x and g(x)= 4/x

OpenStudy (anonymous):

f(x) = 4/x g(x) = 4/x ------------- f(g(x)) = f(4/x) = 4/(4/x) = x g(f(x)) = g(4/x) = 4/(4/x) = x

OpenStudy (anonymous):

THANKS